Lenow Drug Stores and Hall Pharmaceuticals are competitors in the discount drug chain store business. The separate capital structures for Lenow and Hall are presented here.

Lenow Hall
Debt @ 8% $ 300,000 Debt @ 8% $ 600,000
Common stock, $10 par 600,000 Common stock, $10 par 300,000
Total $ 900,000 Total $ 900,000
Common shares 60,000 Common shares 30,000

a. Complete the following table given earnings before interest and taxes of $34,000, $72,000, and $89,000. Assume the tax rate is 10 percent. (Negative amounts should be indicated by parentheses or a minus sign. Round your answers to 2 decimal places.)

What is the relationship between
EBIT Total Assets EBIT/TA % Lenow EPS Hall EPS the EPS of the two firms?
$34,000 $900,000 %
$72,000 $900,000 %
$89,000 $900,000 %

b-1. What is the EBIT/TA rate when the firm's have equal EPS?
  


b-2. What is the cost of debt?
  


b-3. State the relationship between earnings per share and the level of EBIT.
  


c. If the cost of debt went up to 10 percent and all other factors remained equal, what would be the break-even level for EBIT?
Answer #1

ANSWER

a) What is the relationship
EBIT Total Assets EBIT/TA Lenow EPS Hall EPS of the two firms
34000 900000 3.78% $          0.15 $ -0.47 Lenow > Hall
72000 900000 8.00% $          0.72 $ 0.72 Lenow = Hall
89000 900000 9.89% $          0.98 $ 1.23 Lenow < Hall
CALCULATION OF EPS:
LENOW:
EBIT Interest EBT Tax at 10% NI # of shares EPS
34000 24000 10000 1000 9000 60000 $              0.15
72000 24000 48000 4800 43200 60000 $              0.72
89000 24000 65000 6500 58500 60000 $              0.98
HALL:
EBIT Interest EBT Tax at 10% NI # of shares EPS
34000 48000 -14000 0 -14000 30000 $             -0.47
72000 48000 24000 2400 21600 30000 $              0.72
89000 48000 41000 4100 36900 30000 $              1.23
b-1) EBIT/TA rate for equal EPS = 8%
b-2) Cost of debt = 8%
b-3) EPS is unaffected by financial leverage when the pre-tax returns on assets (EBIT/TA) equals the cost of debt.
c) Breakeven EBIT is where the EPS of the two firms are equal.
EPS for Lenow = (E-24000)*90%/60000
EPS for Hall = (E-48000)*90%/30000
where E = the break even EBIT
Equating the two EPS
(E-24000)*90%/60000 = (E-48000)*90%/30000
Solving for E
(E-24000) = 2*(E-48000)
E-24000 = 2*E - 96000
E = 72000
Breakeven EBIT = $24000
CHECK:
Lenow Hall
EBIT 72000 72000
Interest 24000 48000
EBT 48000 24000
Tax at 10% 4800 2400
NI 43200 21600
# of shares 60000 30000
EPS $           0.72 $          0.72

__________________________________

If you have any query or any Explanation please ask me in the comment box, i am here to helps you.please give me positive rating.

*****************THANK YOU*************